# Bounds for Lyapunov exponent of circular light orbits in black holes

### v1

@@ I Introduction
where $\kappa_{H}$ is the surface gravity of the horizon of the black hole and $T_{H}$ its associated temperature.
@@
The Lyapunov exponent associated with unstable circular-photon null geodesics is also related to lensing in the strong deflection limit of black holes Stefanov et al. (2010), to the imaginary part of quasinormal modes in the eikonal limit Cardoso et al. (2009), and to the phase-transition properties of black holes which multiple branches correspond to the distinct phases of the black hole Guo et al. (2022); Kumara et al. (2024); Shukla et al. (2024).
@@
As an application of these new bounds, we derive a constraint for the quasinormal modes. Following the approach in Giataganas et al. (2024), we also present simple bounds for the Lyapunov exponent using the inverse of the entropy of the black hole horizon.
@@ II Background
$p^{\prime}_{r}=$ $-\frac{(e^{-2\delta}\mu)^{\prime}}{2e^{-2\delta}\mu}(\rho+p_{r})+\frac{2}{r}(p_{\bot}-p_{r})\;.w$ (9)
@@ III Bounds on the Lyapunov exponent of circular null geodesics
It results that $\lambda$ is bounded by $\kappa_{\gamma}$ and for the size of the event horizon. In fact, this inequality can be improved for black holes with a traceless energy-momentum tensor. In such cases, Hod recently showed in Hod (2023) that $r_{\gamma}\geq\frac{6}{5}r_{H}$, which implies that $\lambda^{2}r_{H}\leq\frac{5}{6}\kappa_{\gamma}$.
@@
It is worth noting that, recently, in reference Giataganas et al. (2024), the validity of this inequality was observed for the special cases of Schwarzschild and Reissner-Nordström black holes through direct calculation. What we have shown in Eq.(42) is that it remains valid within Einstein’s theory of general relativity, under very natural physical assumptions. In Giataganas et al. (2024), the authors also conjecture new inequalities that have been tested to hold for rotating spacetimes in the Kerr-Newman family (more on this will be discussed below.)
@@
This inequality allows to write a universal bound for the imaginary part of the quasinormal modes in the eikonal limit. In that limit, defining $\omega=\omega_{R}-i\omega_{I}$ as the frequency of the quasinormal modes, we have $\omega_{I}=(n+\frac{1}{2})\lambda$ with $n$ the overtone numberCardoso et al. (2009). Therefore, we obtain the universal upper bound
@@
An associated bound for the real $\omega_{R}$ was derived in Hod (2013).
The bound (45) can be related to a conjecture of Giataganas et al. (2024) relating the Lyapunov exponent to the black hole entropy $S$, which reads Giataganas et al. (2024)
@@ IV Summary
For specific black holes where $|r^{3}p_{r}|$ is bounded, we derived a sharper inequality Eq.(50). As a simple application, we show that these bounds also impose a straightforward constraint on the imaginary part of the quasinormal mode frequencies in the eikonal limit.
It would be interesting to know whether some of the bounds proven here have analogs in other spherically symmetric black holes that do not satisfy the assumptions made in this work, specifically regarding the energy conditions or the requirement of being solutions to the Einstein equations. For instance, a generalized version of (19) was proven in for the $n$-dimensional Einstein-Gauss-Bonnet theory, and in Cvetic et al. (2016) Cvetic, Gibbons and Pope have shown that black holes arising from other gravitational theories also satisfy the bounds found inGallo and Villanueva (2015). Since the Lyapunov bound is, in some sense, related to the upper bound on the location of the photon sphere, it is hoped that more general versions of the bound for $\lambda$ exist for broader classes of metrics or alternative theories. More importantly, it would be desirable to test these inequalities for rotating black holes or to prove their validity in those spacetimes. However, in general, the Einstein equations for rotating spacetimes are more complex, and the null geodesics are typically non-planar. Therefore, further work and potentially different techniques will be required to address this problem.
@@ Acknowledgments
E.G acknowledges financial support from CONICET, SeCyT-UNC. E.G. thanks for the motivating discussions to Frisaldo Hames, Dario Gallardo Mozart, and Guillermo Correa Diaz, while T.M. appreciates discussions with Olivera Mišković, Radouane Gannouji, and Rodrigo Olea.

### v2

@@ I Introduction
where $\kappa_{H}$ is the surface gravity of the horizon of the black hole and $T_{H}$ its associated temperature. For other studies were this bound is satisfied or violated in different spacetimes and gravitational theories we refer to Wang and Chen (2023); Lei et al. (2024); Tang (2024); Addazi et al. (2021); Addazi and Capozziello (2023); Gao et al. (2022); Hashimoto and Sugiura (2023); Yu et al. (2023); Xie et al. (2023); Kan and Gwak (2022); Gwak et al. (2022); Park and Gwak (2024). Alternative methods to study chaos can be found in De Falco and Borrelli (2021a, b).
@@
The Lyapunov exponent associated with unstable circular-photon null geodesics is also related to lensing in the strong deflection limit of black holes Stefanov et al. (2010); Raffaelli (2016), to the imaginary part of quasinormal modes in the eikonal limit Cardoso et al. (2009), and to the phase-transition properties of black holes which multiple branches correspond to the distinct phases of the black hole Guo et al. (2022); Kumara et al. (2024); Shukla et al. (2024).
@@
As an application of these new bounds, we derive a constraint for the quasinormal modes under the assumption of validity of the Wentzel-Kramers-Brillouin (WKB) approximationSchutz and Will (1985) in the eikonal regime. Following the approach in Giataganas et al. (2024), we also present simple bounds for the Lyapunov exponent using the inverse of the entropy of the black hole horizon.
@@ II Background
$p^{\prime}_{r}=$ $-\frac{(e^{-2\delta}\mu)^{\prime}}{2e^{-2\delta}\mu}(\rho+p_{r})+\frac{2}{r}(p_{\bot}-p_{r})\;,$ (9)
@@
For circular null geodesics we also have $E=\pm\frac{\sqrt{e^{-2\delta(r_{\gamma})}\mu(r_{\gamma})}}{r_{\gamma}}L.$
@@
This bound was extended to higher dimensions and various gravitational theories Gallo and Villanueva (2015); Cvetic et al. (2016); Ma and Lu (2020), and to other compact objects Peng (2020); Xavier et al. (2024). Compact objects, including black holes with multiple photon rings, were discussed inCunha et al. (2017); Di Filippo (2024); Xavier et al. (2024); Hod (2018); Guo et al. (2023), and Hod (2023a); Peng (2023) show that extremal black holes generally have external light rings.
@@ III Bounds on the Lyapunov exponent of circular null geodesics
It results that $\lambda$ is bounded by $\kappa_{\gamma}$ and for the size of the event horizon. In fact, this inequality can be improved for black holes with a traceless energy-momentum tensor. In such cases, Hod recently showed in Hod (2023b) that $r_{\gamma}\geq\frac{6}{5}r_{H}$, which implies that $\lambda^{2}r_{H}\leq\frac{5}{6}\kappa_{\gamma}$.
@@
It is worth noting that, recently, in reference Giataganas et al. (2024), the validity of this inequality was observed for the special cases of Schwarzschild and Reissner-Nordström black holes through direct calculation. What we have shown in Eq.(42) is that it remains valid within Einstein’s theory of general relativity, under very natural physical assumptions. In Giataganas et al. (2024), the authors also conjecture new inequalities that have been tested to hold for rotating spacetimes in the Kerr-Newman family (more on this will be discussed below). Note also, that a relation between the Lyapunov exponent and the Unruh temperature associated to a scalar field thermalization in the vicinity of an effective metric near to the photon sphere was also observed in Raffaelli (2022).
@@
This inequality allows to write a bound for the imaginary part of the quasinormal modes in the eikonal limit, assuming that the WKB approximation can be usedSchutz and Will (1985); Konoplya et al. (2019). In that limit, defining $\omega=\omega_{R}-i\omega_{I}$ as the frequency of the quasinormal modes, we have $\omega_{I}=(n+\frac{1}{2})\lambda$ with $n$ the overtone numberCardoso et al. (2009). Therefore, for those quasinormal modes we obtain the upper bound
@@
An associated bound for the real $\omega_{R}$ was derived in Hod (2013). However, this correspondence between null geodesics and quasinormal modes should be taken with caution. First, as shown by Konoplya in Konoplya and Stuchlík (2017), this correspondence can only be guaranteed in spacetimes where the effective potential associated to the quasinormal modes has a single peak, decaying at the event horizon and infinity, and it is limited to test fields and not for gravitational and other non-minimally coupled field. As shown in Konoplya and Stuchlík (2017); Konoplya (2023); Bolokhov (2024), this might not be the case in many alternative theories or even in asymptotically flat spacetimes. Additionally, the issue may be that the WKB approximation is valid but does not produce the complete spectrum of quasinormal modesKonoplya (2023). Therefore, the previous result should be considered valid only for those modes where the correspondence can be guaranteed, but it should not be taken as universal. As discussed in Konoplya and Stuchlík (2017), the same applies to the bounds for the real part of these modes as those found in Gallo and Villanueva (2015); Hod (2013).
Referring back to the bound (45), it is also related to a conjecture of Giataganas et al. (2024) relating the Lyapunov exponent to the black hole entropy $S$, which reads Giataganas et al. (2024)
@@
To conclude this section, we would like to relate some of the bounds found here to an alternative approach for studying circular null geodesics and their stability. This method, developed in Cunha et al. (2022) (see also Qiao (2022); Qiao and Li (2022)), is based on a two-dimensional Riemannian optical metricGordon (1923); Gibbons and Werner (2008); Crisnejo and Gallo (2018), which has the property of sharing the spatial orbits of null geodesics of the spacetime as its own geodesics. More precisely, for a static and spherically symmetric spacetime of the form
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{h(r)}+r^{2}d\Omega^{2},$ (51)
one assigns an optical metric:
$dt^{2}=\frac{1}{f(r)}\left(\frac{dr^{2}}{h(r)}+r^{2}d\Omega^{2}\right).$ (52)
The stability of circular null geodesics at $r_{\gamma}$ follows from the sign of the Gaussian curvature $\mathcal{K}$ associated with these orbits, which at the position $r_{\gamma}$ is given byCunha et al. (2022)
$\mathcal{K}(r_{\gamma})=\frac{h(r_{\gamma})}{2}\left(f^{\prime\prime}(r_{\gamma})-\frac{f^{\prime}(r_{\gamma})}{r_{\gamma}}\right).$ (53)
If $\mathcal{K}<0$, the orbits are unstable. After straightforward calculations, it can be verified using the expression for (28) for the Lyapunov exponent, and $N(r_{\gamma})=0$, that
$\mathcal{K}(r_{\gamma})=-\lambda^{2}.$ (54)
This relation can be easily understood by taking into account that the geodesic deviation equation in optical space for the deviation vector field $y^{a}$ along the circular null geodesic at $r_{\gamma}$ is a system of second order linear differential equations which linearly depend on the Gaussian curvatureCasetti et al. (1996); do Carmo (2016).
Thus, taking into account the inequality (45), we can observe that at the position $r_{\gamma}$ of the innermost unstable null circular orbit of a black hole with event horizon at $r_{H}$, the Gaussian curvature of the associated optical metric is bounded from below and above as
$-\frac{1}{3r_{H}^{2}}\leq\mathcal{K}<0,$ (55)
that is, we have found that the optical space cannot have arbitrarily negative curvature at the unstable photon sphere.
@@ IV Summary
For specific black holes where $|r^{3}p_{r}|$ decreases monotonically, we derived a sharper inequality Eq.(50). As a simple application, we show that these bounds also impose a straightforward constraint on the imaginary part of the quasinormal mode in the regime where the correspondence between circular null geodesics and quasinormal modes is valid.
It would be interesting to know whether some of the bounds proven here have analogs in other spherically symmetric black holes that do not satisfy the assumptions made in this work, specifically regarding the energy conditions or the requirement of being solutions to the Einstein equations. For instance, a generalized version of (19) was proven in for the $n$-dimensional Einstein-Gauss-Bonnet theory (see also Ghosh et al. (2023)), and in Cvetic et al. (2016) Cvetic, Gibbons and Pope have shown that black holes arising from other gravitational theories also satisfy the bounds found inGallo and Villanueva (2015). Since the Lyapunov bound is, in some sense, related to the upper bound on the location of the photon sphere, it is hoped that more general versions of the bound for $\lambda$ exist for broader classes of metrics or alternative theories. More importantly, it would be desirable to test these inequalities for rotating black holes or to prove their validity in those spacetimes. However, in general, the Einstein equations for rotating spacetimes are more complex, and the null geodesics are typically non-planar. Therefore, further work and potentially different techniques will be required to address this problem.
@@ Acknowledgments
E.G. acknowledges financial support from CONICET and SeCyT-UNC. We also thank R. Konoplya for bringing attention to relevant works on the WKB approximation to quasinormal modes. E.G. thanks Frisaldo Hames, Dario Gallardo Mozart, and Guillermo Correas Diaz for motivating discussions, while T.M. appreciates discussions with Olivera Mišković, Radouane Gannouji, and Rodrigo Olea.